# BOP-D: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities

### v1

@@ Abstract
Currently, 6D pose estimation methods are benchmarked on datasets that consider, for their ground truth annotations, visual ambiguities as only related to global object symmetries. However, as previously observed [26], visual ambiguities can also happen depending on the viewpoint or the presence of occluding objects, when disambiguating parts become hidden. The visual ambiguities are therefore actually different across images. We thus first propose an automatic method to re-annotate those datasets with a 6D pose distribution specific to each image, taking into account the visibility of the object surface in the image to correctly determine the visual ambiguities. Given this improved ground truth, we re-evaluate the state-of-the-art methods and show this greatly modify the ranking of these methods. Our annotations also allow us to benchmark recent methods able to estimate a pose distribution on real images for the first time. We will make our annotations for the T-LESS dataset and our code publicly available.
Figure 1: We provide for the first time 6D pose annotations in the form of a per-image distribution. Current annotations in BOP [15] datasets are given as a single pose, shown here as a circle in the SO(3) representations. T-LESS [11] also provides a symmetry pattern leaving the 3D object model invariant, from which a distribution can be computed (the colored points in the SO(3) representations). Such distribution however does not cover many cases [26]: For the object in this example, when only its core is visible (Case 1), the pose is fully ambiguous and should be represented by a continuous distribution in SO(3). When the sides of the head are visible (Case 2), there are still ambiguities and the distribution is made of 6 clusters. When the hole is visible (Case 3), there are no ambiguities, and the pose distribution should be concentrated around one pose. Our method annotates the T-LESS dataset with per-image distributions, taking into account the partial occlusions and allowing us to evaluate a predicted pose properly. We show that considering these distributions for evaluating state-of-the-art methods results in a significant change of ranking for the BOP challenge.
@@ 1 Introduction
Visual 6D pose estimation of an object consists in determining the 3D position and orientation of this object with respect to the camera that explain the observed image. It is a key task in many application domains, such as robotics (e.g. grasping or manipulation), augmented reality, industrial quality control, etc. In such contexts, datasets of images with pose ground-truths are of primary importance, since they can be used to train learning based methods, but first and foremost to benchmark the performances of proposed approaches. The annotation accuracy of reference benchmarks, such as BOP [39] for 6D pose estimation, is thus crucial since it influences the research directions.
@@
In this paper, we first propose a method that automatically estimates pose distributions for images that take into account the ambiguities due to the object symmetries but also to the partial occlusions. Our method exploits the current annotations as a starting point, the ground truth visibility masks of the objects, and their 3D models. While we demonstrate it on T-LESS, our method is fully automatic and can be applied to other datasets.
We also re-evaluate classical metrics with these annotations to evaluate the performance of pose estimation methods that return a single pose estimate per image. This yields a strongly modified ranking of current challenger methods.
@@
    A novel automatic method computing a multi-modal 6D pose distributions ground truth from a unique ground truth pose and a proposal set of object symmetries;
@@ 2 Related Work
While the state of the art mostly focused on estimating a single pose [4, 35, 13, 21, 38, 41, 6, 37], some recent works focus on estimating the complete set of solutions [26, 33, 30, 20, 27, 16, 7, 17].
As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are required, even for methods that returns a single pose per image.
@@
For single pose estimation, the accuracy is measured through the deviation of the object surface points when transformed by the estimate and by the ground truth pose (registration error) [10]. Depending on the sensor used—RGB or RGB-D—and the targeted application (e.g. robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface. In case of multi-valued ground-truths due to a symmetrical object or occlusion-induced symmetries, the accuracy corresponds to the minimal accuracy with respect to the set of solutions [10, 44, 40]. In practice, the most commonly used accuracy measures are :
@@
    Average Distance metric (ADD), and its variations for symmetrical objects (ADD-S [44], ADD-H [40]), target to measure the mean Euclidean error between the surface points of the object transformed with the estimated pose and the ground truth pose, but have been replaced by MSSD now.
@@
    Visual Surface Discrepancy (VSD) [12] measures the ratio of pixels for which the discrepancy between their value in the distance map computed with the ground truth pose and their values in the distance map computed with the pose to evaluate. VSD is more expensive to compute than MSSD and MSPD, and requires a depth image. It is now omitted for new tasks as stated by the BOP organizers on their website [^1] https://bop.felk.cvut.cz/method_info/447/.
@@
Regarding the evaluation of multi-modal 6D pose distribution estimation methods, the accuracy is usually measured as a pose error, meaning the minimal deviation between the inferred rotation and translation and the corresponding nearest ground truth poses [30, 7, 18, 36].
@@
However, not all multi-modal 6D pose estimation methods actually provide a full probability distribution over the whole pose space. Instead, some methods output a set of poses corresponding to some local maxima of the underlying probability distribution [17]. For such methods, performances can be evaluated through the Precision/Recall of the multi-valuated estimation. A pose estimate is then considered as correct with respect to one ground truth pose if its distance in pose space do not exceed a threshold $\delta$. The Precision for a given image is then defined as the ratio of the number of estimated poses whose distance have at least one ground truth pose that do not exceed the threshold $\delta$ over the total number of estimated poses. The Recall for an object image is then defined as the ratio between the number of ground truth pose whose distance have at least one estimated pose that do not exceed the threshold $\delta$ over the total number of ground truth poses. Similarly to the spread and log likelihood, these precision and recall are related to the multi-modal pose distribution.
@@ 3 Method
Depending on the geometry $M$ of the object and the set ${\cal V}(M)$ of its visible points, set ${\cal T}$ can be limited to the identity transform in the case of a non=ambiguous pose or can include multiple transformations in addition to the identity. As introduced in [10], ${\cal T}$ can be defined as partial $\epsilon$-symmetries of ${\cal V}(M)$:
@@
where $d$ is the distance between the surfaces and $\epsilon$ the surface deviation tolerance.
If the set ${\cal V}(M)$ of visible points is available, the symmetry pattern could in principle be approximated by sampling the rigid transformation space SE(3) and assessing whether the sample belongs to the $\epsilon$-symmetries. However, since sampling a 6D space becomes quickly computationally expensive, such annotation process was never used at our knowledge.
In the following sections, we introduce solutions to reduce the computational cost (Section 3.1), how to introduce some tolerance in the evaluation of the $\epsilon$-symmetries (Section 3.2), and how to extend the annotation to textured object (Section 3.3).
@@ 3.1 Speeding up the computation of $\epsilon$-sym
First, by definition of the $\epsilon$-symmetries, for any subsets ${\cal V}_{1}$ and ${\cal V}_{2}$ of visible points of a 3D model $M$, we have:
$\epsilon\text{-sym}({\cal V}_{1}\cup{\cal V}_{2},M)=\epsilon\text{-sym}({\cal V}_{1},M)\cap\epsilon\text{-sym}({\cal V}_{2},M)\>.$ (2)
Therefore, for any subset ${\cal V}=\{v_{j}\}$ of visible points of a 3D object $M$, we have:
$\epsilon\text{-sym}({\cal V},M)=\bigcap_{j}\epsilon\text{-sym}(v_{j},M)\>,$ (3)
where $\epsilon\text{-sym}(v_{j},M)$ is the symmetry pattern when the point $v_{j}$ is the only visible point, which we name elementary symmetry pattern of $v_{j}$.
@@ 3.2 Robust symmetries patterns intersection
A strict intersection would only keep the $T_{i}$ with the maximum value of $\text{H}(T_{i})$ which is represented by $\text{H}(Id)$ as the identity of SE(3) and is always in the symmetry pattern. Our soft intersection tolerates a threshold and keeps all the $T_{i}$’s such that $\text{H}(T_{i})>\text{H}(Id)-\tau$, where $\tau$ represents the minimum size (expressed in number of 3D points) of the symmetry-breaking element.
@@ 4 Evaluating 6D Pose Distribution Predictions
Currently, single pose estimation methods and distribution pose estimation methods are evaluated on different datasets (BOP challenge for single poses, SYMSOL for 6D pose distributions) with different metrics (registration error for single pose estimation, pose error for 6D pose distribution).
With our annotation method that provides the full set of 6D pose solutions even on real data, it becomes possible to evaluate both tasks on the same dataset. However, it would be beneficial if both tasks can be evaluated with comparable metrics.
We therefore propose to keep the evaluation process of single pose method unchanged, but on a more accurate ground truth, and to extend it to pose distributions evaluation.
First, whereas error pose is commonly used in 6D pose distribution evaluation process [30, 18, 7], we propose to use the registration error. The latter takes the object 3D shape into account, making the accuracy more meaningful for most of the application. It is also the type of error used by the current gold standard benchmark for single pose [39]. Typically, we suggest to use the Maximal Surface Distance (MSD) and Maximum Projective Distance (MPD) metrics.
Second, we propose to keep the metrics of Precision and Recall, but we adapt them to measure Precision over the whole estimated distribution of poses instead of a unique pose, and Recall over the whole set of solutions instead of considering a pose to be found if one of its multiple solutions was found.
@@
where Est and GT are respectively the estimated and the ground truth distributions to compare, and $p(\text{Est}_{i})$ represents the probability associated to the $i$-th element of the estimated distribution  [^2] If the method output solutions without assigning a probability to each pose, a uniform probability over the solutions is used. , $\mathbf{d}$ is the chosen registration distance, and $\tau_{\mathbf{d}}$ the associated threshold.
@@ 5 Experiments
Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset. Indeed, among the datasets of the BOP challenge [15], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [5] and Homebrewed [19] feature symmetrical objects, with occlusion-induced symmetries, but do not share publicly their ground-truths. HOPE [40], IC-BIN [3] and YCB-V [44] have symmetrical shapes, but texture information disambiguate them completely.
The first experiment focus on a qualitative evaluation of our annotation method (Section 5.2). The second on evaluate the impact of our re-annotation of T-LESS Dataset on the performance evaluation of the best performer methods of the BOP benchmark [15] on single pose estimation  (the Multiple instances of Multiple known objects (MiMo) task of BOP) (Section 5.3). The last experiment evaluate for the first time methods of 6D pose distribution estimation on a non-synthetic dataset with proper annotations and metrics (Section 5.4).
@@ 5.1 Implementation details
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points.
@@ 5.2 Qualitative validation of the method
For the qualitative evaluation of our annotating method, we illustrate the results obtained on some images of the T-LESS dataset in Figure 3.
@@
As expected, we can observe the coherence between the distribution and the visibility of the disambiguating elements. Since this dataset does not contain any image with visual ambiguity for the translational part of the pose, the ground truth translation is always reduced to a single 3D point.
More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
@@
Figure 3: Visualizations of our ground truth. We display SE(3) ground truth distribution for scenes 12, 16 and 17. Circle on orientation diagram represent the unique ground truth pose that is provided as input to our method. Colors are used to link object instances to their distributions. More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
BOP [39] BOP-D (our annotations)
Methods Modality $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ Rank $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ Rank
T-LESS [11] Challenge: pose estimation
gpose2023 [45] RGB-D 92.1 94.6 93.4 1 86.5 89.1 87.8 1
gdrnppv2-rgbd-pbrreal [41, 25] RGB-D 90.2 92.6 91.4 2 84.8 87.2 86.0 3 ( $\downarrow$1)
gpose2023-pbr [41, 25] RGB-D 90.3 92.6 91.4 3 84.7 87.2 86.0 2 ( $\uparrow$1)
modalocclusion-rgbd RGB-D 89.7 92.9 91.3 4 83.2 86.4 84.8 4
gdrbpp v1.5 [41, 25] RGB-D 88.7 91.5 90.1 5 58.1 61.9 60.0 14 ( $\downarrow$9)
gdrnpp v1.4 [41, 25] RGB-D 88.4 90.9 89.6 6 60.6 63.7 62.2 10 ( $\downarrow$4)
gdrnpp v1.3 [41, 25] RGB-D 87.1 90.3 88.7 7 57.5 61.2 59.3 16 ( $\downarrow$9)
zebraposesat-effnetb4 [37] RGB-D 87.3 90.2 88.7 8 60.2 63.4 61.8 11 ( $\downarrow$3)
gdrnpp-pbr-rgbd-mmodel v1.2 [41, 25] RGB-D 87.2 90.1 88.7 9 60.8 64.2 62.5 9
hccepose RGB 81.4 94.8 88.1 10 53.1 64.5 58.8 18 ( $\downarrow$8)
hipose-cvpr24 [24] RGB-D 86.6 88.8 87.7 11 58.0 60.8 59.4 15 ( $\downarrow$4)
cosypose [21] RGB 83.6 90.7 87.2 12 57.2 64.0 60.6 13 ( $\downarrow$1)
pfa RGB-D 85.5 87.7 86.6 13 59.2 62.0 60.6 12 ( $\uparrow$1)
hccepose RGB 79.2 93.2 86.2 14 51.4 63.1 57.2 20 ( $\downarrow$6)
zebraposesa [37] RGB 79.1 93.0 86.1 15 51.8 63.4 57.6 19 ( $\downarrow$4)
gdrnppde v1.1 [41] RGB-D 78.4 92.2 85.3 16 69.9 82.9 76.4 7 ( $\uparrow$9)
leroy-fuseocclu-rgb RGB-D 78.0 91.7 84.8 17 70.4 83.3 76.9 6 ( $\uparrow$11)
gdpnpp v1.0 [41, 25] RGB-D 83.4 85.6 84.5 18 57.5 60.2 58.9 17 ( $\uparrow$1)
surfemb [6] RGB-D 82.9 85.9 84.4 19 75.8 78.7 77.2 5 ( $\uparrow$14)
poseio RGB-D 81.7 86.0 83.8 20 69.3 73.6 71.4 8 ( $\uparrow$12)

### v2

# BOP-Distrib: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities

@@ Abstract
6D pose estimation aims at determining the pose of the object that best explains the camera observation. The unique solution for a non-symmetrical object can turn into a multi-modal pose distribution for a symmetrical object or when occlusions of symmetry-breaking elements happen, depending on the viewpoint. Currently, 6D pose estimation methods are benchmarked on datasets that consider, for their ground truth annotations, visual ambiguities as only related to global object symmetries, whereas they should be defined per-image to account for the camera viewpoint. We thus first propose an automatic method to re-annotate those datasets with a 6D pose distribution specific to each image, taking into account the visibility of the object surface in the image to correctly determine the visual ambiguities. Second, given this improved ground truth, we re-evaluate the state-of-the-art single pose methods and show that this greatly modifies the ranking of these methods. Third, as some recent works focus on estimating the complete set of solutions, we derive a precision/recall formulation to evaluate them against our image-wise distribution ground truth, making it the first benchmark for pose distribution methods on real images. We will make our annotations for the T-LESS dataset and our code publicly available.
Figure 1: We provide for the first time 6D pose annotations in the form of a per-image object pose distribution. Current annotations in BOP [17] datasets are given as a single pose, shown here as a circle in the SO(3) representations. T-LESS [13] also provides a symmetry pattern leaving the 3D object model invariant, from which a distribution can be computed (the colored points in the SO(3) representations). Such distribution however does not cover many cases [29]: For the object in this example, when only its core is visible (Case 1), the pose is fully ambiguous and should be represented by a continuous distribution in SO(3). When the sides of the head are visible (Case 2), there are still ambiguities and the distribution is made of 6 modes. When the hole is visible (Case 3), there are no ambiguities, and the pose distribution should be concentrated around one pose. Our method annotates the T-LESS dataset with per-image distributions, taking into account the partial occlusions and allowing us to evaluate a predicted pose properly. We show that considering these distributions for evaluating state-of-the-art methods results in a significant change of ranking for the BOP challenge. Such ground truth distributions also become a key asset when it comes to evaluating pose distribution estimation methods [9, 19]. With appropriate metrics, we demonstrate the first quantitative evaluation of pose distribution methods on real images, as an extension to single pose methods.
@@ 1 Introduction
Visual 6D pose estimation of an object consists in determining the 3D position and orientation of this object with respect to the camera that explains the observed image. It is a key task in many application domains, such as robotics (e.g. grasping or manipulation), augmented reality, industrial quality control, etc. In such contexts, datasets of images with ground-truth poses are of primary importance, since they can be used to train learning based methods, but first and foremost to benchmark the performances of proposed approaches. The annotation accuracy of reference benchmarks, such as BOP [42] for 6D pose estimation, is thus crucial since it influences the research directions.
@@
In this paper, we first propose a method that automatically annotates pose distributions for images that take into account the ambiguities due to the object symmetries but also to the partial occlusions. Our method exploits the current annotations as a starting point, the ground truth visibility masks of the objects, and their 3D models. While we demonstrate it on T-LESS [13], our method is fully automatic and can be applied to other datasets as demonstrated on YCB-V [47].
We also re-evaluate classical metrics with these annotations to assess the performance of pose estimation methods that return a single pose estimate per image. This yields a strongly modified ranking of current challenger methods.
@@
    A novel automatic method computing a multi-modal 6D pose distribution ground truth from a unique ground truth pose and a proposal set of object symmetries;
@@ 2 Related Work
While the state of the art mostly focused on estimating a single pose [6, 38, 15, 23, 41, 44, 8, 40], some recent works focus on estimating the complete set of solutions [29, 36, 33, 22, 30, 18, 9, 19]. As we argue below, to properly evaluate the performance of a method, ground truth annotations including the complete set of solutions for each test image, and metrics that take into account the multiplicity of the solutions, are required, even for methods that return a single pose per image.
@@
For single pose estimation, the accuracy is measured through the deviation of the object surface points when transformed by the estimate and by the ground truth pose (registration error) [12]. Depending on the sensor used—RGB or RGB-D—and the targeted application (e.g. robotics or Augmented Reality), deviations can be measured in 3D space (3D distance) or in the image space (reprojection error), and the registration error can be considered as the mean or maximal error over the object surface. In case of multi-valued ground-truths due to a symmetrical object or occlusion-induced symmetries, the accuracy corresponds to the minimal error with respect to the set of solutions [12, 47, 43]. In practice, the most commonly used accuracy measures are :
@@
    Average Distance metric (ADD), and variations for symmetrical objects (ADD-S [47], ADD-H [43]), target to measure mean Euclidean error between surface points of the object transformed with the estimated pose and the ground truth pose, but are replaced by MSSD now.
@@
    Visual Surface Discrepancy (VSD) [14] measures misalignment over the visible surface of the object model. VSD is more expensive to compute than MSSD and MSPD, and requires a depth image. It is now omitted for new tasks as stated by BOP organizers [^1] https://bop.felk.cvut.cz/challenges/.
@@
Regarding the evaluation of multi-modal 6D pose distribution estimation methods, the accuracy is usually measured as a pose error, meaning the minimal deviation between the inferred rotation and translation and the corresponding nearest ground truth pose [33, 9, 20, 39].
@@
However, not all multi-modal 6D pose estimation methods actually provide a full probability distribution over the whole pose space. Instead, some methods output a set of poses corresponding to some local maxima of the underlying probability distribution [19]. For such methods, performances can be evaluated through the Precision/Recall of the multi-valuated estimation. A pose estimate is then considered as correct with respect to one ground truth pose if its distance in pose space does not exceed a threshold $\delta$. The Precision for a given image is then defined as the ratio of the number of estimated poses whose distance have at least one ground truth pose that does not exceed the threshold $\delta$ over the total number of estimated poses. The Recall for an object image is then defined as the ratio between the number of ground truth poses whose distance have at least one estimated pose that does not exceed the threshold $\delta$ over the total number of ground truth poses. Similarly to the spread and log likelihood, these precision and recall are related to the multi-modal pose distribution.
@@ 3 Method
Depending on the geometry $M$ of the object and the set ${\cal V}(M)$ of its visible points, set ${\cal T}$ can be limited to the identity transform in the case of a non-ambiguous pose or can include multiple transformations in addition to the identity. As introduced in [12], ${\cal T}$ can be defined as partial $\epsilon$-symmetries of ${\cal V}(M)$:
@@
where $d$ measures the mean deviation between the points of ${\cal V}(M)$ and $M$, $\epsilon$ is the maximal surface deviation tolerance and * represents the application of the transformation $T_{i}$ onto the point cloud ${\cal V}(M)$.
If the set ${\cal V}(M)$ of visible points is available, the symmetry pattern could in principle be approximated by sampling the rigid transformation space SE(3) and assessing whether the sample belongs to the $\epsilon$-symmetries. However, since sampling a 6D space becomes quickly computationally expensive, such annotation process was never used to our knowledge.
In the following sections, we introduce solutions to reduce the computational cost (Section 3.1), how to introduce some tolerance in the evaluation of the $\epsilon$-symmetries (Section 3.2), and how to extend the annotation to textured objects (Section 3.3).
@@ 3.1 Speeding up the computation of $\epsilon$-sym
First, by definition of the $\epsilon$-symmetries, for any subsets $V_{1}\subset{\cal V}(M)$ and $V_{2}\subset{\cal V}(M)$, we have:
$\epsilon\text{-sym}(V_{1}\cup V_{2},M)=\epsilon\text{-sym}(V_{1},M)\cap\epsilon\text{-sym}(V_{2},M)\>.$ (2)
Therefore, we can reformulate Equation 1 as:
$\epsilon\text{-sym}({\cal V}(M),M)=\bigcap_{v_{j}\in{\cal V}(M)}\epsilon\text{-sym}(v_{j},M)\>,$ (3)
where $v_{j}$ is a 3D point and $\epsilon\text{-sym}(v_{j},M)$ is the symmetry pattern when the point $v_{j}$ is the only visible point, which we name elementary symmetry pattern of $v_{j}$.
@@ 3.2 Robust symmetries patterns intersection
A strict intersection would only keep the $T_{i}$ with the maximum value of $\text{H}(T_{i})$ which is represented by $\text{H}(Id)$ as the identity of SE(3) and is always in the symmetry pattern. Our soft intersection tolerates a threshold and keeps all the $T_{i}$ such that $\text{H}(T_{i})>\text{H}(Id)-\tau$, where $\tau$ represents the minimum size (expressed in number of 3D points) of the symmetry-breaking elements.
@@ 4 Evaluating 6D Pose Distribution Predictions
Currently, single pose estimation methods and distribution pose estimation methods are evaluated on different datasets (BOP challenge for single poses, SYMSOL for 6D pose distributions) with different metrics (registration error for single pose estimation, pose error for 6D pose distribution). With our annotation method that provides the full set of 6D pose solutions even on real data, it becomes possible to evaluate both tasks on the same dataset. However, it would be beneficial if both tasks can be evaluated with comparable metrics.
We therefore propose to keep the evaluation process of single pose method unchanged, but on a more accurate ground truth, and to extend it to pose distributions evaluation. First, whereas pose error is commonly used in 6D pose distribution evaluation process [33, 20, 9], we propose to use the registration error. The latter takes the object 3D shape into account, making the accuracy more meaningful for most of the applications. It is also the type of error used by the current gold standard benchmark for single pose [42]. Typically, we suggest to use the Maximal Surface Distance (MSD) and Maximum Projective Distance (MPD) metrics. Second, we propose to keep the metrics of Precision and Recall, but we adapt them to measure Precision over the whole estimated distribution of poses instead of a unique pose, and Recall over the whole set of solutions instead of considering a pose to be found if one of its multiple solutions was found.
@@
where Est and GT are respectively the estimated and the ground truth distributions to compare, and $p(\text{Est}_{i})$ represents the probability associated to the $i$-th element of the estimated distribution  [^2] If the method outputs solutions without assigning a probability to each pose, such as [19], a uniform probability over the solutions is used., $\mathbf{d}$ is the chosen registration distance, and $\tau_{\mathbf{d}}$ the associated threshold.
@@ 5 Experiments
Since our annotation process and metrics are related to objects with intrinsic or occlusion-induced symmetries, we perform our evaluations on the T-LESS dataset. Indeed, among the datasets of the BOP challenge [17], T-LESS appears to be the only dataset to exhibit symmetrical objects, with occlusion-induced symmetries, real data and public ground-truth annotations: ITODD [7] and Homebrewed [21] feature symmetrical objects, with occlusion-induced symmetries, but do not publicly share their ground-truths. HOPE [43] and IC-BIN [5] have symmetrical shapes, but texture information disambiguates them completely.
The first experiment focus on a qualitative evaluation of our annotation method (Section 5.2). The second evaluates the impact of our re-annotation of T-LESS Dataset on the performance evaluation of the best performer methods of the BOP benchmark [17] on single pose estimation  (the Multiple instances of Multiple known objects (MiMo) task of BOP) (Section 5.3). The last experiment evaluates for the first time methods of 6D pose distribution estimation on a non-synthetic dataset with proper annotations and metrics (Section 5.4).
@@ 5.1 Implementation details
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 2.5$\times$2.5mm2.
@@ 5.2 Qualitative validation of the method
We illustrate qualitatively the results obtained on some images of the T-LESS dataset in Figure 3. We can observe the coherence between the distribution and the visibility of the disambiguating elements. Since this dataset does not contain any image with visual ambiguity for the translational part of the pose, the ground truth translation is always reduced to a single 3D point. More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
@@
Figure 3: Visualizations of our ground truth. We display SE(3) ground truth distributions for scene 16 of T-LESS [13]. Circle on orientation diagram represents the unique ground truth pose provided as input to our method. Colors are used to link object instances to their distributions. More visualizations are provided in the supplementary material to convince the reader of the quality of our annotations.
BOP [42] (object-wise annotations) BOP-Distrib (our image-wise annotations)
BOP Contenders $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ Rank $\mathbf{MSSD}$ $\mathbf{MSPD}$ $\mathbf{Mean}$ $\mathbf{Loss}$ Rank
T-LESS [13] BOP Challenge 2023 [17]: pose estimation of unseen objects
foundpose [37] 55.0 62.3 58.6 1 45.5 52.1 48.8 -9.8 3( $\downarrow$-2)
gigaposemegapose-5-hypothesis [35, 24] 54.3 62.0 58.1 2 45.3 52.3 48.8 -9.3 2
gigaposemegapose-1-hypothesis [35, 24] 51.5 59.2 55.3 3 38.5 45.1 41.8 -13.5 9( $\downarrow$-6)
genflow-multihypo16-rgb [32] 50.9 57.9 54.4 4 45.9 52.4 49.1 -5.3 1( $\uparrow$3)
genflow-multihypo-rgb [32] 50.4 57.3 53.9 5 44.9 51.4 48.1 -5.8 4( $\uparrow$1)
genflow-multihypo16 [32] 52.8 54.8 53.8 6 46.7 48.6 47.7 -6.1 5( $\uparrow$1)
genflow-multihypo [32] 51.6 53.6 52.6 7 45.2 47.2 46.2 -6.4 6( $\uparrow$1)
foundpose [37] 49.1 55.6 52.4 8 37.7 43.4 40.6 -11.8 11( $\downarrow$-3)
foundpose [37] 49.1 55.6 52.3 9 37.7 43.4 40.5 -11.8 12( $\downarrow$-3)
cnos-fastsammegapose-multihyp-10 [34, 24] 48.5 55.9 52.2 10 40.6 47.1 43.8 -8.4 7( $\uparrow$3)
cnos-fastsammegapose-multihyp [34, 24] 48.4 55.6 52.0 11 40.5 47.1 43.8 -8.2 8( $\uparrow$3)
RGB cnos-fastsammegapose-multihyp-teaserpp [34, 24] 48.4 51.3 49.9 12 40.3 43.0 41.7 -8.2 10( $\uparrow$2)
foundationpose [46] 62.9 64.1 63.5 1 56.7 58.0 57.4 -6.1 1
gigaposegenflowkabsch [35] 56.3 58.6 57.5 2 49.4 51.7 50.6 -6.9 2
sam6d [26] 52.9 54.5 53.7 3 39.7 41.2 40.4 -13.3 3
sam6d [26] 52.1 53.5 52.8 4 38.3 39.7 39.0 -13.8 5( $\downarrow$-1)
sam6d-fastsam [26] 49.9 51.1 50.5 5 37.9 39.1 38.5 -12.0 6( $\downarrow$-1)
sam6d-cnosmask [26] 50.0 50.9 50.4 6 39.3 40.3 39.8 -10.6 4( $\uparrow$2)
sam6d-cnosfastsam [26] 49.3 50.5 49.9 7 37.2 38.4 37.8 -12.1 7
sam6d-zeropose [26] 44.4 45.6 45.0 8 33.5 34.8 34.2 -10.8 8
RGB-D zeropose-multi-hypo-refinement-defaultseg [4] 41.5 42.0 41.8 9 29.4 30.1 29.7 -12.1 9
T-LESS [13] BOP Challenge 2022 [42]: pose estimation
hccepose-2024-pbr 82.9 95.8 89.4 1 53.8 65.2 59.5 -29.9 5( $\downarrow$-4)
hcceposeefficientnet-b4-default-2d-bbox 81.4 94.8 88.1 2 53.1 64.5 58.8 -29.3 6( $\downarrow$-4)
cosypose-eccv20-syntreal-8views [23] 83.6 90.7 87.2 3 57.2 64.0 60.6 -26.6 4( $\downarrow$-1)
hccepose-default-2d-bbox 79.2 93.2 86.2 4 51.4 63.1 57.2 -29.0 8( $\downarrow$-4)
gdrnppdet-pbrrealgenflow-multihypo-rgb [44, 28] 78.4 92.2 85.3 5 69.9 82.9 76.4 -8.9 2( $\uparrow$3)
leroy-fuseocclu-rgb 78.0 91.7 84.8 6 70.4 83.3 76.9 -7.9 1( $\uparrow$5)
gpose2023-rgb [48] 76.6 92.9 84.7 7 50.2 64.0 57.1 -27.6 9( $\downarrow$-2)
gdrnpp-pbr-rgb-mmodel [44, 28] 76.3 92.4 84.4 8 50.0 63.6 56.8 -27.6 10( $\downarrow$-2)
gdrnpp-pbrreal-rgb-mmodel [44, 28] 76.0 91.3 83.6 9 48.8 61.7 55.2 -28.4 13( $\downarrow$-4)
cosypose-eccv20-syntreal-4views [23] 79.5 86.4 83.0 10 52.7 59.3 56.0 -27.0 12( $\downarrow$-2)
mrc-net [25] 78.5 87.1 82.8 11 52.3 60.3 56.3 -26.5 11
gdrnppdet-pbrrealmegapose-multihyp [44, 28] 74.0 87.4 80.7 12 61.6 73.8 67.7 -13.0 3( $\uparrow$9)
RGB sc6d [3] 72.1 85.3 78.7 13 52.6 64.1 58.3 -20.4 7( $\uparrow$6)
gpose2023 [48] 92.1 94.6 93.4 1 86.5 89.1 87.8 -5.6 1
gdrnppv2-rgbd-pbrreal [44, 28] 90.2 92.6 91.4 2 84.8 87.2 86.0 -5.4 3( $\downarrow$-1)
gpose2023-pbr [48] 90.3 92.6 91.4 3 84.7 87.2 86.0 -5.4 2( $\uparrow$1)
modalocclusion-rgbd 89.7 92.9 91.3 4 83.2 86.4 84.8 -6.5 4
hcceposebf-2024-ref 89.0 91.7 90.3 5 71.7 74.7 73.2 -17.1 6( $\downarrow$-1)
gdrbpp-pbrreal-rgbd-mmodel-fast v1.4 [44, 28] 88.7 91.5 90.1 6 58.1 61.9 60.0 -30.1 13( $\downarrow$-7)
gdrnpp-pbrreal-rgbd-mmodel v1.3 [44, 28] 88.4 90.9 89.6 7 60.6 63.7 62.2 -27.4 10( $\downarrow$-3)
gdrnpp-pbrreal-rgbd-smodel v1.2 [44, 28] 87.1 90.3 88.7 8 57.5 61.2 59.3 -29.4 15( $\downarrow$-7)
zebraposesat-effnetb4-refineddefdet-2023 [40] 87.3 90.2 88.7 9 60.2 63.4 61.8 -26.9 11( $\downarrow$-2)
gdrnpp-pbr-rgbd-mmodel v1.1 [44, 28] 87.2 90.1 88.7 10 60.8 64.2 62.5 -26.2 9( $\uparrow$1)
hipose-cvpr24 [27] 86.6 88.8 87.7 11 58.0 60.8 59.4 -28.3 14( $\downarrow$-3)
defaultdetection-pfa-mixpbr-rgb-d 85.5 87.7 86.6 12 59.2 62.0 60.6 -26.0 12
zebraposesat-effnetb4defdet-2023 [40] 79.1 93.0 86.1 13 51.8 63.4 57.6 -28.5 17( $\downarrow$-4)
zebraposesat-effnetb4pbr-only-defdet-2023 77.7 92.0 84.9 14 50.7 62.8 56.8 -28.1 18( $\downarrow$-4)
gdpnpp-pbrreal-rgbd-mmodel-officialdet v1.0 [44, 28] 83.4 85.6 84.5 15 57.5 60.2 58.9 -25.6 16( $\downarrow$-1)
surfemb-pbr-rgbd-lin [8] 82.9 85.9 84.4 16 75.8 78.7 77.2 -7.2 5( $\uparrow$11)
RGB-D poseio 81.7 86.0 83.8 17 69.3 73.6 71.4 -12.4 8( $\uparrow$9)
YCB-V [47] BOP Challenge 2022 [42]: pose estimation
cosypose-eccv20-syntreal-8views [23] 88.5 88.0 88.2 1 83.3 82.4 82.9 -5.3 2( $\downarrow$-1)
cosypose-eccv20-syntreal-4views [23] 86.9 86.6 86.8 2 82.1 81.3 81.7 -5.1 4( $\downarrow$-2)
mrpe-pbrreal-rgb-smodel 85.7 87.3 86.5 3 83.1 84.4 83.7 -2.8 1( $\uparrow$2)
RGB magic 85.9 85.4 85.6 4 82.9 82.3 82.6 -3.0 3( $\uparrow$1)
gdrnppv2-rgbd-pbrreal [44, 28] 96.3 92.7 94.5 1 94.2 90.2 92.2 -2.3 1
gpose2023 [48] 96.2 92.6 94.4 2 94.2 90.2 92.2 -2.2 2
gdrnpp-pbrreal-rgbd-mmodel [44, 28] 95.7 91.9 93.8 3 90.1 86.2 88.1 -5.7 3
RGB-D dsgc6d 93.0 90.1 91.5 4 87.5 85.2 86.4 -5.1 4